# Retroreflective surface optimisation for optical cavities with custom mirror profiles

### v1

@@ V Example: Multiple emitters in cavities
The case of the single emitter at the centre of the cavity indicates that mirror surface optimisation can find modes that circumvent traditional limitations of Gaussian beams and consequently improve performance. In that case, the improvement was derived from finding a mode that could fit the shape of the mirror more efficiently than a Gaussian mode, and thus use all the available space on the end mirrors to achieve a tight focus. We now discuss applications involving multiple emitters in a single optical cavity, to ascertain whether improvements could be found for these use cases.
@@ V.1 Coupling of two axial emitters
We will start with the simplest case of two identical emitters equally spaced to either side of the centre of the cavity (see Fig. 2a). The enforced axial symmetry of our chosen solutions guarantees that the internal cooperativity is the same on both emitters and therefore we can use the gradient-based optimisation discussed in Sec. II.
@@
We first discuss our qualitative expectations of behaviour with spherical mirrors. If the separation $s$ of the emitters is small, the task of focussing strongly on both emitters is identical to the task of focussing to the centre of the cavity. This is the exact case covered in Sec. IV, where the performance is limited by the diameter $D$ of the mirrors, which places an upper limit on the divergence of a low loss cavity mode. However, beyond a certain emitter separation $s$, a cavity mode optimised for central intensity will have diverged significantly at the positions of both emitters, and therefore we would prefer to reduce the mode divergence down from the upper limit. In this regime, the performance is limited by the separation of the emitters, not the diameter of the mirrors. We expect the transition between these two regimes to occur at a separation $s_{\mathrm{crit}}$ for which both limitations are saturated simultaneously. In this case, the best Gaussian mode for minimising beam size on both emitters has clipping loss equal to $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$. This leads to
$s_{\mathrm{crit}}=\frac{\lambda L^{2}}{\pi D^{2}}\left\{\log(2)-\log\left(\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)\right\},$ (15)
@@
derived in App. D. Assuming $s>s_{\mathrm{crit}}$, the best Gaussian mode is limited by the requirement that it does not diverge strongly between the centre of the cavity and the emitters. As derived in App. D, the optimum radius of curvature and internal cooperativity are
$R=\frac{L}{2}\left(1+\frac{s^{2}}{L^{2}}\right),\quad C_{\mathrm{int}}^{\mathrm{min}}=\frac{6\lambda}{\pi s\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}}.$ (16)
@@
We again study the ability of retroreflective optimisation to improve the performance in this scenario by performing example calculations, as displayed in Fig. 3. We again select 136 free optimisation parameters in a larger calculation basis of 171 states to converge over the range of geometries. We choose an example for which $s>s_{\mathrm{crit}}$, and find that retroreflective optimisation does significantly increase performance (Fig. 3a). The optimised mirror shape has an inner and outer radius of curvature, with an abrupt transition between, and interference patterns around that transition (Fig. 3b). The central fit curvature of the left (right) mirror and outer fit curvature of the right (left) mirror have a geometric focus on the left (right) ion. This produces a mode which focusses strongly on both emitters (Fig. 3d). When scanning different geometries Fig. 3e), we see that, where $s<s_{\mathrm{crit}}$, the best spherical and optimised $C_{\mathrm{int}}^{\mathrm{min}}$ are fairly constant. This is because the limitation to performance is the ability of the geometry to produce a strongly focussing mode, which does not depend on the separation of the emitters, and is roughly determined by the ratio $D/L$, which is constant in the data. However, where $s>s_{\mathrm{crit}}$, the performance of the cavity with spherical mirrors is now limited by the separation of the emitters, and so decreases as $s$ increases. The optimised mode retains high $C_{\mathrm{int}}^{\mathrm{min}}$, because it can focus effectively on both emitters. This can lead to improvements exceeding an order of magnitude in $C_{\mathrm{int}}^{\mathrm{min}}$ over a wide range of geometries Fig. 3f).
@@
Figure 3: Example and summary of improvements in $C_{\mathrm{int}}^{\mathrm{min}}$ achievable with retroreflective optimisation for two emitters placed on the axis of the cavity, symmetrically about the central point (see Fig. 2a). The example cavity (panels a-d) has length $2.8\text{\,}\mathrm{mm}$, mirror diameter $0.56\text{\,}\mathrm{mm}$, resonant wavelength $854\text{\,}\mathrm{nm}$, emitter separation $1.12\text{\,}\mathrm{mm}$, and non-clipping internal loss 20ppm. a) $C_{\mathrm{int}}^{\mathrm{min}}$ across the two emitters for (black line) two identical spherical mirrors as a function of mirror curvature $1/R$ with (dashed black) expected optimum $C_{\mathrm{int}}^{\mathrm{min}}$ and $1/R$ from Eq. 16, and (horizontal blue line) for retroreflective optimised mirrors. b) Surface profile of the (black) best spherical mirror and (blue) optimised mirror, with (red) residuals overlaid. The optimised surface is fit to a profile with an abrupt transition (dashed green vertical line) between an inner (dotted green line) and outer (dashed green line) spherical profile. c) Mode intensities (linear scale in top row, log scale in bottom row) for the best spherical mirror cavity (black) and the optimised cavity (blue) in i) a transverse plane containing one of the emitters and ii) the mirror transverse plane, with vertical lines marking the edges of the mirror. d) Mode intensities of the i) best spherical mirror cavity, ii) retroreflective-optimised cavity in an xzxz cross-section of the mode ($y=0$) with (cyan crosses) the two emitter positions. In ii) The surface fit from b) is indicated through green arrows to show the scale of the central (outer) curvature $R_{c}$ ($R_{o}$) and green dashed lines to show the focal points of the central curvature from the transition radius, and the outer curvature from the mirror radius. e) $C_{\mathrm{int}}^{\mathrm{min}}$ across the two emitters for spherical mirrors and optimised mirrors. The white dashed diagonal line depicts $s_{\mathrm{crit}}$, and the vertical line delineates $D<D_{\mathrm{crit}}$ (left) from $D>D_{\mathrm{crit}}$ (right). f) Improvement in the $C_{\mathrm{int}}^{\mathrm{min}}$ value from retroreflective optimisation. No data shown where $D<D_{\mathrm{crit}}$. The green crosses in ei), eii), and f) show the configuration used for panels a)-d).
@@ V.2 Emitters on an axial line
Example data is shown in Fig. 4, where we optimise to couple the cavity mode to an array of $N_{e}=10$ emitters. Here, we see three distinct regimes in terms of performance level and improvement, determined by the relation of the length of the array ($L_{e}$), and the separation of the emitters ($s=L_{e}/(N_{e}-1)$) to the critical separation $s_{\mathrm{crit}}$ (see Eq. 15). In the short array case ($L_{e}<s_{\mathrm{crit}}$) the array is so short that the optimum fundamental mode is found by focusing the Gaussian beam as tightly as possible to the centre of the cavity. In this case, the spherical mirror and optimised mirror $C_{\mathrm{int}}^{\mathrm{min}}$ values are, again, roughly independent of the array geometry, with improvement factors around the level found for the case of Sec. IV. In the second ‘intermediate’ regime $s<s_{\mathrm{crit}}<L_{e}$, the array length necessitates that the optimum spherical mirror is less concentric than the short array regime, but because $s>s_{\mathrm{crit}}$, the optimised mode cannot focus on each emitter individually without diverging too strongly to be contained by the mirrors of diameter $D$. Empirically, we see that the $C_{\mathrm{int}}^{\mathrm{min}}$ of the spherical and optimised mirror cavities drops with increasing array length in this regime. In the final regime, $s<s_{\mathrm{crit}}$, and the optimised mirrors are now able to focus the mode on each emitter individually (Fig. 4IId). This intensity pattern does look like a standing wave, but it actually arises from interference of transverse modes with a periodicity much bigger than the standing wave. In this regime, the $C_{\mathrm{int}}^{\mathrm{min}}$ of the spherical mirror cavity continues to fall with increasing array length, but the optimised performance plateaus, allowing for larger improvement factors. This metric could not be optimised with a gradient-based optimiser, so we instead used a simplex optimisation strategy, which is likely the reason for the increased numerical noise on the data. The data taken for specific examples again used 136 free parameters and 171 basis states, but to make the optimisation less numerically heavy, the scan data used 81 free parameters and 111 basis states, which was found sufficient for performance convergence.
@@
Figure 4: Example and summary of improving $C_{\mathrm{int}}^{\mathrm{min}}$ across an axial emitter array using retroreflective optimisation. The example cavity has length $2.8\text{\,}\mathrm{mm}$, mirror diameter $0.56\text{\,}\mathrm{mm}$, resonant wavelength $854\text{\,}\mathrm{nm}$, non-clipping internal loss 20ppm, and $N_{e}=10$ equispaced emitters. The short (long) array of panels I (II) has length $0.28\text{\,}\mathrm{mm}$ ($1.12\text{\,}\mathrm{mm}$). a) Achievable $C_{\mathrm{int}}^{\mathrm{min}}$ for (black line) two identical spherical mirrors as a function of mirror curvature and (horizontal blue line) retroreflective optimised mirrors. The dashed black lines show the expected optimum $C_{\mathrm{int}}^{\mathrm{min}}$ and corresponding radius of curvature from Eq. 16, but where the $s$ in that expression is taken be $L_{e}$, the separation of the two furthest emitters. b) The surface profile of the best spherical mirror (black) and the optimised mirror (blue), with the residuals overlaid (red). c) Mode intensities of the i) best spherical mirror cavity ii) retroreflective-optimised cavity in an xzxz cross-section of the mode ($y=0$). The cyan crosses mark the emitter positions. d) The intensity of the optimised mode along the cavity axis. The vertical lines indicate the positions of the emitters. e) The maximum $C_{\mathrm{int}}^{\mathrm{min}}$ across the array achievable with ei) spherical mirrors and eii) optimised mirrors. The lower white dashed diagonal line depicts $L_{e}=s_{\mathrm{crit}}$, the upper white diagonal line $s=s_{\mathrm{crit}}$, and the white vertical line $D=D_{\mathrm{crit}}$. f) The improvement in $C_{\mathrm{int}}^{\mathrm{min}}$ from choosing the optimised surface. No data shown where $D<D_{\mathrm{crit}}$. The green crosses in ei), eii), and f) show the configuration used for panels a)-d), where the cross label ‘I’ or ‘II’ corresponds to the example case.
@@
VI Conclusion

### v2

@@ IV.3 Possible improvements in performance
IV.4 Non axially-centred emitter
The case where the emitter is in the centre of the cavity is the most natural for achieving strong coupling, but wider experimental factors may impose dissimilar geometric constraints on the mirrors, or place the emitter away from the centre of the cavity. In this case, we expect the optimised cavity to have an asymmetric mode, with the mirrors on the left and right having different profiles. Retroreflective optimisation can also be used in this case
Example data is shown in Fig. 2, which studies a cavity where the emitter is moved towards the right mirror, and additionally where the right mirror is larger than the left. To achieve the strongest focus at the emitter, the cavity with spherical mirror must find the best compromise between two factors. On one hand, the Gaussian mode would ideally focus on the emitter so that the highest intensity of the mode is at the pertinent location. However, in this case, it is the left mirror, which is both smaller and further from the emitter than the right mirror, that limits the mode divergence. By moving the focus towards the left mirror, more strongly focussing modes can fit inside the cavity without suffering overwhelming clipping loss. Ultimately, this balance means that the best spherical mirrors produce a Gaussian mode that focuses slightly closer to the centre of the cavity than the emitter (Fig. 2a). The retroreflective optimised mode again fits precisely onto the left mirror, whereas the optimised intensity behaves less dramatically on the right mirror, because clipping here does not limit performance (Fig. 2c and d). The optimised mode is also slightly asymmetric about the basis focal position, something which the Gaussian mode cannot do, although the factor 2.75 improvement in $C_{\mathrm{int}}$ is broadly in line with those found in the symmetric case.
[…]
IV.5 Transversely-displaced emitter
Having seen that mirror shaping can improve the coupling cooperativity of a single emitter in the centre of a symmetric cavity, and that improvements can also be found when the cavity geometry is not symmetric about the axis, we now discuss whether the cooperativity can be improved if the emitter is displaced from the centre of the cavity in a transverse direction. As this breaks the azimuthal symmetry, we consider a cavity with just one transverse dimension, using the Hermite-Gauss modes as our basis states.
We simulate two example cases (Fig. 3), where the top row shows the emitter displaced a small amount from the axis, and the bottom row the emitter displaced close to in line with the edges of the mirrors. In the former case, we see that the best spherical mirrors focus just slightly more centrally in the cavity than the position of the emitter in order to focus most tightly (Fig. 3Ia). The optimised mode again uses the space on the mirror more efficiently to achieve a tighter focus (1.65 times improvement in $C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}$, see Fig. 3Ib), requiring small but precise modifications to the surface profile (Fig. 3Ic).
[…]
In the case of large displacement, the physics is somewhat different. Most notably, the spherical mirrors that lead to the highest performance are all found for beam family waists corresponding precisely to transverse degeneracies. The waist values can be predicted as
$w_{\mathrm{res}}^{q,p}=\sqrt{\frac{\lambda L}{2\pi\tan\left(\frac{\pi q}{2p}\right)}},$ (16)
where $q$ is the number of free spectral ranges between degenerate modes, and $p$ is the increase in the transverse mode index $n$ between degenerate modes. These degeneracies allow the cavity modes to mix in order to avoid clipping on the mirror edge, as can be seen visually in Fig. 3IIb) (see [41] for further discussion on transverse mode mixing). The optimised mode uses asymmetry precisely to create a tightly focused mode with strong divergence towards the cavity centre, but little divergence away from it Fig. 3IIb) (bottom). This improves $C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}$ by 2.45 times from the best circular mirror case, which itself had precisely selected mirror curvature to satisfy the mode degeneracy condition.
@@ V Example: Multiple emitters in cavities
The case of the single emitter indicates that mirror surface optimisation can find modes that circumvent traditional limitations of Gaussian beams and consequently improve performance. In that case, the improvement was derived from finding a mode that could fit the shape of the mirror more efficiently than a Gaussian mode, and thus use all the available space on the end mirrors to achieve a tight focus. We now discuss applications involving multiple emitters in a single optical cavity, to ascertain whether improvements could be found for these use cases.
@@ V.1 Coupling of two axial emitters
We will start with the simplest case of two identical emitters equally spaced to either side of the centre of the radially symmetric cavity (see Fig. 4a). The enforced axial symmetry of our chosen solutions guarantees that the internal cooperativity is the same on both emitters and therefore we can use the gradient-based optimisation discussed in Sec. II.
@@
We first discuss our qualitative expectations of behaviour with spherical mirrors. If the separation $s$ of the emitters is small, the task of focussing strongly on both emitters is identical to the task of focussing to the centre of the cavity. This is the exact case covered in Sec. IV, where the performance is limited by the diameter $D$ of the mirrors, which places an upper limit on the divergence of a low loss cavity mode. However, beyond a certain emitter separation $s$, a cavity mode optimised for central intensity will have diverged significantly at the positions of both emitters, and therefore we would prefer to reduce the mode divergence down from the upper limit. In this regime, the performance is limited by the separation of the emitters, not the diameter of the mirrors. We expect the transition between these two regimes to occur at a separation $s_{\mathrm{crit}}$ for which both limitations are saturated simultaneously. In this case, the best Gaussian mode for minimising beam size on both emitters has clipping loss equal to $\mathcal{L}_{\mathrm{abs}}$. This leads to
$s_{\mathrm{crit}}=\frac{\lambda L^{2}}{\pi D^{2}}\left\{\log(2)-\log\left(\mathcal{L}_{\mathrm{abs}}\right)\right\},$ (17)
@@
derived in App. E. Assuming $s>s_{\mathrm{crit}}$, the best Gaussian mode is limited by the requirement that it does not diverge strongly between the centre of the cavity and the emitters. As derived in App. E, the optimum radius of curvature and internal cooperativity are
$R=\frac{L}{2}\left(1+\frac{s^{2}}{L^{2}}\right),\quad C_{\mathrm{int}}^{\mathrm{min}}=\frac{6\lambda}{\pi s\mathcal{L}_{\mathrm{abs}}}.$ (18)
@@
We again study the ability of retroreflective optimisation to improve the performance in this scenario by performing example calculations, as displayed in Fig. 5. We again select 136 free optimisation parameters in a larger calculation basis of 171 Laguerre-Gauss basis states to converge over the range of geometries. We choose an example for which $s>s_{\mathrm{crit}}$, and find that retroreflective optimisation does significantly increase performance (Fig. 5a). The optimised mirror shape has an inner and outer radius of curvature, with an abrupt transition between, and interference patterns around that transition (Fig. 5b). The central fit curvature of the left (right) mirror and outer fit curvature of the right (left) mirror have a geometric focus on the left (right) ion. This produces a mode which focusses strongly on both emitters (Fig. 5d). When scanning different geometries Fig. 5e), we see that, where $s<s_{\mathrm{crit}}$, the best spherical and optimised $C_{\mathrm{int}}^{\mathrm{min}}$ are fairly constant. This is because the limitation to performance is the ability of the geometry to produce a strongly focussing mode, which does not depend on the separation of the emitters, and is roughly determined by the ratio $D/L$, which is constant in the data. However, where $s>s_{\mathrm{crit}}$, the performance of the cavity with spherical mirrors is now limited by the separation of the emitters, and so decreases as $s$ increases. The optimised mode retains high $C_{\mathrm{int}}^{\mathrm{min}}$, because it can focus effectively on both emitters. This can lead to improvements exceeding an order of magnitude in $C_{\mathrm{int}}^{\mathrm{min}}$ over a wide range of geometries (Fig. 5f).
@@
Figure 5: Example and summary of improvements in $C_{\mathrm{int}}^{\mathrm{min}}$ achievable with retroreflective optimisation for two emitters placed on the axis of the cavity, symmetrically about the central point (see Fig. 4a). The example cavity (panels a-d) has length $2.8\text{\,}\mathrm{mm}$, mirror diameter $0.56\text{\,}\mathrm{mm}$, resonant wavelength $854\text{\,}\mathrm{nm}$, emitter separation $1.12\text{\,}\mathrm{mm}$, and non-clipping internal loss 20ppm. a) $C_{\mathrm{int}}^{\mathrm{min}}$ across the two emitters for (black line) two identical spherical mirrors as a function of mirror curvature $1/R$ with (dashed black) expected optimum $C_{\mathrm{int}}^{\mathrm{min}}$ and $1/R$ from Eq. (18), and (horizontal blue line) for retroreflective optimised mirrors. b) Surface profile of the (black) best spherical mirror and (blue) optimised mirror, with (red) residuals overlaid. The optimised surface is fit to a profile with an abrupt transition (dashed green vertical line) between an inner (dotted green line) and outer (dashed green line) spherical profile. c) Mode intensities (linear scale in top row, log-10 scale in bottom row with gridlines indicating decades) for the best spherical mirror cavity (black) and the optimised cavity (blue) in i) a transverse plane containing one of the emitters and ii) the mirror transverse plane, with vertical lines marking the edges of the mirror. d) Mode intensities of the i) best spherical mirror cavity, ii) retroreflective-optimised cavity in an xzxz cross-section of the mode ($y=0$) with (cyan crosses) the two emitter positions. In ii) The surface fit from b) is indicated through green arrows to show the scale of the central (outer) curvature $R_{c}$ ($R_{o}$) and green dashed lines to show the focal points of the central curvature from the transition radius, and the outer curvature from the mirror radius. e) $C_{\mathrm{int}}^{\mathrm{min}}$ across the two emitters for spherical mirrors and optimised mirrors. The white dashed diagonal line depicts $s_{\mathrm{crit}}$, and the vertical line delineates $D<D_{\mathrm{crit}}$ (left) from $D>D_{\mathrm{crit}}$ (right). f) Improvement in the $C_{\mathrm{int}}^{\mathrm{min}}$ value from retroreflective optimisation. No data shown where $D<D_{\mathrm{crit}}$. The green crosses in ei), eii), and f) show the configuration used for panels a)-d).
@@ V.2 Emitters on an axial line
Example data is shown in Fig. 6, where we optimise to couple the cavity mode to an array of $N_{e}=10$ emitters. Here, we see three distinct regimes in terms of performance level and improvement, determined by the relation of the length of the array ($L_{e}$), and the separation of the emitters ($s=L_{e}/(N_{e}-1)$) to the critical separation $s_{\mathrm{crit}}$ (see Eq. (17)). In the short array case ($L_{e}<s_{\mathrm{crit}}$) the array is so short that the optimum fundamental mode is found by focusing the Gaussian beam as tightly as possible to the centre of the cavity. In this case, the spherical mirror and optimised mirror $C_{\mathrm{int}}^{\mathrm{min}}$ values are, again, roughly independent of the array geometry, with improvement factors around the level found for the case of Sec. IV. In the second ‘intermediate’ regime $s<s_{\mathrm{crit}}<L_{e}$, the array length necessitates that the optimum spherical mirror is less concentric than the short array regime, but because $s>s_{\mathrm{crit}}$, the optimised mode cannot focus on each emitter individually without diverging too strongly to be contained by the mirrors of diameter $D$. Empirically, we see that the $C_{\mathrm{int}}^{\mathrm{min}}$ of the spherical and optimised mirror cavities drops with increasing array length in this regime. In the final regime, $s<s_{\mathrm{crit}}$, and the optimised mirrors are now able to focus the mode on each emitter individually (Fig. 6IId). This intensity pattern does look like a standing wave, but it actually arises from interference of transverse modes with a periodicity much bigger than the standing wave. In this regime, the $C_{\mathrm{int}}^{\mathrm{min}}$ of the spherical mirror cavity continues to fall with increasing array length, but the optimised performance plateaus, allowing for larger improvement factors. This metric could not be optimised with a gradient-based optimiser, so we instead used a simplex optimisation strategy, which is likely the reason for the increased numerical noise on the data.
@@
Figure 6: Example and summary of improving $C_{\mathrm{int}}^{\mathrm{min}}$ across an axial emitter array using retroreflective optimisation. The example cavity has length $2.8\text{\,}\mathrm{mm}$, mirror diameter $0.56\text{\,}\mathrm{mm}$, resonant wavelength $854\text{\,}\mathrm{nm}$, non-clipping internal loss 20ppm, and $N_{e}=10$ equispaced emitters. The short (long) array of panels I (II) has length $0.28\text{\,}\mathrm{mm}$ ($1.12\text{\,}\mathrm{mm}$). a) Achievable $C_{\mathrm{int}}^{\mathrm{min}}$ for (black line) two identical spherical mirrors as a function of mirror curvature and (horizontal blue line) retroreflective optimised mirrors. The dashed black lines show the expected optimum $C_{\mathrm{int}}^{\mathrm{min}}$ and corresponding radius of curvature from Eq. (18), but where the $s$ in that expression is taken be $L_{e}$, the separation of the two furthest emitters. b) The surface profile of the best spherical mirror (black) and the optimised mirror (blue), with the residuals overlaid (red). c) Mode intensities of the i) best spherical mirror cavity ii) retroreflective-optimised cavity in an xzxz cross-section of the mode ($y=0$). The cyan crosses mark the emitter positions. d) The intensity of the optimised mode along the cavity axis. The vertical lines indicate the positions of the emitters. e) The maximum $C_{\mathrm{int}}^{\mathrm{min}}$ across the array achievable with ei) spherical mirrors and eii) optimised mirrors. The lower white dashed diagonal line depicts $L_{e}=s_{\mathrm{crit}}$, the upper white diagonal line $s=s_{\mathrm{crit}}$, and the white vertical line $D=D_{\mathrm{crit}}$. f) The improvement in $C_{\mathrm{int}}^{\mathrm{min}}$ from choosing the optimised surface. No data shown where $D<D_{\mathrm{crit}}$. The green crosses in ei), eii), and f) show the configuration used for panels a)-d), where the cross label ‘I’ or ‘II’ corresponds to the example case.
@@
VI Prospect for manufacture
The retroreflective optimisation method approximately finds the best possible cavity performance under geometrical constraints for arbitrary mirror shaping. This method is thus well-suited to finding the performance limits imposed by the laws of physics and the corresponding mirror shapes. Perhaps most valuably, it shows how significantly mirror shaping could improve performance in the target application without assuming a limited set of surface parameterisations.
The method does not, however, guarantee that the optimised surfaces are easy to manufacture, and the optimised surfaces (see Figs. 1, 2, 3, 5, and 6) typically harbour high-spatial frequency features. To find the precision of the spatial shaping required to make these mirrors, we take example optimised surfaces, remove frequency components above a cutoff shaping scale with a fourth-order Butterworth filter, and find the performance of cavities with the resulting surfaces. Example results are shown in Fig. 7 which indicate (row ii) very low performance if the cutoff scale is too low and near-optimised performance for high spatial resolution. Across the examples, the transition between these low and high performance regions occurs at a shaping scale of approximately
$(1/\lambda_{s})_{\mathrm{crit}}=\frac{D}{\lambda L},$ (19)
which is the spatial frequency that would diffract a beam from parallel to the cavity axis to parallel with the line joining the edges of opposite mirrors (i.e. the diffraction allowed by the cavity solid angle). However, the sharp features of the mirror that couples the cavity mode to two emitters (Fig. 7 column c) requires shaping slightly more precise than this.
[…]
These required shaping scales, though still much larger than the smallest beam diameters of focused ion beam milling [29], are very demanding. We would therefore recommend that retroreflective optimisation is used as part of a wider method to find promising mirror shapes for manufacture, rather than designing the surfaces itself. In this approach, the retroreflective optimisation method identifies applications for which mirror shaping holds promise to significantly outperform cavities with spherical mirrors, inspires simpler mirror designs that slightly compromise performance for improved manufacturability, and contextualises the performance of these simpler designs against the ultimate physical limits of the geometry. For a full example of this procedure applied to plano-concave cavities see Hughes and Horak [74].
VII Conclusion